# Importance of accounting for student identities and intersectionality for creating equitable and inclusive physics learning environments

### v1

@@ III.5 Supports and Role Models
In this section, we share who or what has supported and sustained these women and kept them on their path during their undergraduate physics journey. This support may include role models, i.e., people who they looked up to. In the previous themes, such as Family providing academic support, we discussed how some of the women already highlighted how their family provided support by helping them with their homework. For many of these women, their family members were also role models, especially if they had careers in science. For example, Cathy shares that her sister and her parents were some of her role models and provided foundations and inspiration for going into a science major. Many women noted that their professors were their role models. Some of these professors were also their academic or research advisors, thus serving multiple roles and having multiple contexts for providing support.
@@
Chloe shares that one of her professors is really inspiring because of a shared identity. Chloe considers herself to be an introvert and her professor is also quiet, so she feels inspired by her: “my astronomy professor is a big inspiration…she’s the first real astronomer that I’ve met and particularly that she is kind of a quieter person and most of the professors that I’ve met, you ask them a question and then they talk very confidently for hours and she’s not like that…so meeting somebody who’s not like that but is still very successful in the field, yeah, that is very inspiring.” This type of role model is valuable as the stereotypical image of a physicist is usually either social and outgoing (like those portrayed in the media) or awkward and standoffish. Seeing someone who shares similar characteristics is helpful for someone like Chloe, who does not have the stereotypical personality of a physicist, to see herself as capable of being successful as one.
@@
Francis explains that having female professors to look up to and who act as role models help her feel like she belongs. “I think there’s quite a few female professors here and I feel like having that role model of a woman who is in physics is definitely nice.” Again, this shared identity can provide comfort for students and like in Francis’s case, help her feel like she belongs.
Another important source of support Chloe mentions is mentoring. The mentoring program is part of a gender inclusive STEM club where upperclassmen partner with lowerclassmen. Chloe specifies that mentoring program is all girls and the upperclassmen “lead [the lowerclassmen] through the classes, and so my big sib[ling] was really helpful and showing me that I can apply for these research and these internships and to pursue these opportunities, yeah and just seeing people who are older and are currently doing what I am aspiring to do in the future is very inspiring.” This mentoring program connected her with more experienced students and particularly her mentor who guided her through opportunities such as research programs and internships.
III.6 Suggestions to improve experiences
@@
In this section, we present personal advice that has helped these women persist in physics and suggestions for instructors they think would help improve their experiences.
III.6.1 Personal Advice: The ”Band-Aid” Approach
@@
We note that although this personal advice may have helped a few of these women, it may not be useful for others. We also labeled this personal advice as The ”Band-Aid” Approach” because although this may be genuine advice from these interviewed women in physics, it only scratches the surface of the issues that traditionally marginalized students face within physics environments. In particular, this advice is more oriented towards personal action as opposed to more systemic change in the physics culture, which is discussed in a later sub-theme.
@@
Cathy has served on an undergraduate committee for physics majors in her program. She says that this opportunity has been beneficial for her and that “having a lot of leadership roles available has been helpful for me because you can get involved in making whatever changes you feel would be helpful and having the department listen to the students…” Through her role on this committee, she was also able to host events for her peers: “I felt like that gave me a lot more power to just host the events that I’ve been wishing we had started up, like study groups for classes and stuff, and started up a physics GRE program, study program so I liked that once I was involved, I could then kind of make the events that I thought would be helpful so I’m glad I did that.” It appears that this may be personal advice based specifically upon the experience with undergraduate major committee that helped Cathy. We note that not every student may have this opportunity or may feel safe being in these types of roles. However, being involved in different types of departmental committees and activities based upon interest may be valuable for students.
@@
Ruby provides some advice that has helped her persist, which includes having a growth mindset: “I think what has helped me the most is learning that it’s okay to not understand everything and that most people don’t even know really anything but it’s about just practicing and trying and also just learning as much as you can. And I think that advice has really helped me in that feeling support, even when I’m not going to be even close to the smartest person in that room. [It] has really helped me.” Having growth mindset may in fact be useful for students to adopt. However, it is especially effective when instructors and the physics culture encourages and reflects a growth mindset as well so that the onus is not only on the student, like Ruby, to adopt a growth mindset in order to excel in physics. In particular, when instructors have growth mindset about their students’ potential and believe that all students can do well in physics, it can help students develop growth mindset and believe that they can become good at physics by working hard, working smart and taking advantage of resources.
@@
Francis shares that she thinks that believing in yourself would help future women in physics “believe in yourself and just because other people, especially men, are talking more and acting like they know more, it’s not necessarily true and you know, [you] are just as intelligent as them.” Thus, Francis’ advice based upon her personal experiences is to “believe in yourself” and although this is a good thing to do, it may be difficult for women in physics if they do not have the support and it does not appear to them that other people believe in them (as in the case of Paulina). This advice may be viewed by some as putting onus on the student to change to fit in to survive and make it through the physics culture. Having instructors believe in their students is necessary for traditionally marginalized students to truly believe in themselves with regard to excelling in physics. These positive changes can be heralded with changing the actual culture of physics that can be usually intimidating for many students.
III.6.2 Advice for instructors
@@
Avoid forming opinions about students based on background: Karina shares some insight on how instructors should not make assumptions about students based on their background: “I think [it is important] for professors to know that [they] try their best to avoid forming opinions on students. In that way, especially a small school like [my college], where you might have professors for multiple years, you might come back to that professor that you’ve had your first semester and I think that’s important, that’s something that’s important for professors to know.” This is especially important, as Karina emphasizes because the college is small and students may have the same professor for multiple courses throughout their program. Besides potentially having the same instructor for multiple courses, these instructors also serve as research advisors and mentors, so it is especially important to have good perceptions of (and relationships with) students. In addition, having preconceived notions about students may unintentionally lead to instructors stereotyping students or making generalizations, particularly about their ability.
@@
Addressing and dismantling biases against women: Ruby shares that it is important for instructors to explicitly communicate in their actions and words that women are just as capable as men in terms of their physics ability: “making it a very, very clear fact that…men are no better at physics than women. Just because there are more men in the history of physics does not mean that men in your classroom are going to be better at it than you are because of your gender, because that has no influence on how capable you are [of] learning the subject, even if people tell you otherwise.” Changing the physics culture to reflect beliefs about all students as equally capable of excelling in physics is key. Having instructors emphasize this and continue to do so not only through verbal affirmations but via their actions as well would be necessary to reflect how deeply they support this tenet.
@@
Chloe shares that she feels like she has to prove herself to people because they do not think that she is capable. She thinks that people make assumptions that men are capable: “I feel like as a woman, I have to prove myself for people to look at me and go, oh, you’re smart and dedicated and you like science, I have to prove that first. It’s not implicitly assumed and nobody assumes that when they first meet me, well, I think, for many men in my classes it’s the opposite, where it’s like you initially assume that this person is really smart and gets really good grades just because they’re male in science.” She believes increasing the number of women in physics programs could help reduce these biases: “I think the only way to fix it…[is] if many other girls are also doing really well, I think it’s a self perpetuating cycle that as more women get involved in science, the more everybody gets, realizes that we’re on the same playing field.” Again, Chloe feels like she has to prove herself, so having instructors dismantle these stereotypes and biases about who belongs in physics and can excel in it through their actions can be very impactful.
@@
Reassurance about lack of representation of women in physics: Ruby shares that she is aware of the historical lack of representation of women in physics but she would appreciate her department to acknowledge that: “I think it’d be great to address the fact that there are not many historical women in physics, but that does not mean that there cannot be in the future, women who make an impact in physics and I think making that clear is so important because I’ve never really heard that before like I haven’t really ever been told …just because there aren’t a lot of women that did it before you, doesn’t mean you can’t do it, but it’d be so reassuring and helpful to hear…I think that that would be beneficial to sort of creating that comfortable environment and creating that confidence.” Thus, Ruby emphasizes that acknowledging this part of physics history would not only educate departmental members but also students would be aware of this going forward in their major. Following this acknowledgement, reassurance would be helpful so that women are not worried about majoring in physics and the department feels supportive to them.
@@
Create a better representation of women in physics: Esme shares that having more representation of women in physics would be beneficial: “seeing more women be the forefront, be the face of physics, because usually when we think of physics, we think of the men…which you know they’ve done a lot for us, but you know, seeing more women just in their discoveries is something that probably can be helpful.” Thus, based upon Esme’s suggestion, highlighting more female physicists such as in lectures, seminars, newsletters, etc. can provide support for women. This can show them the achievements of female physicists who have made major contributions to physics historically and in the present, which can potentially improve their sense of belonging and persistence in physics.
@@
Provide positive recognition: Ruby shares that some reassurance would be helpful especially when she does not feel confident in herself. For example, not doing well on an exam would be an opportunity to reassure students: “more reassurance would be nice in a sense. Sometimes it’s scary and sometimes it makes you feel like am I in this and not able to do this?, and a lot of the times, my confidence is definitely shaken up by a test grade or something along those lines…I think, on the part of the school and professors [to be] making it more clear that it’s okay to not always get the right answer…and that doesn’t mean you’re stupid and that doesn’t mean that you shouldn’t do what you’re passionate about…” In Ruby’s opinion, instructors should make the learning environment safe so that students do not feel bad about getting a wrong answer. Reframing mindsets about wrong answers as an opportunity to learn would be useful in addition to instructors reassuring students that they do belong in physics and that it is okay to not do well on an exam.
@@
Improving office hours: Ruby provides several suggestions for instructors to improve their classrooms, one of them being improving the perception and purpose of office hours: “I think there’s not as safe environment as there should be for getting things wrong or for not knowing things and I think…making a more accepting environment and comfortable environment to make mistakes and to not always know the answer…a lot of the time I think not knowing the answer honestly would help a lot, so that people can feel confident, even when they don’t know everything, because there’s this big, I think, problem with the culture, people are terrified to even try to answer a question in the fear of getting it wrong, and I think that shouldn’t be the case, the best way to learn is by making mistakes and by getting wrong because that’s the only way you can grow. So I think changing that part of the culture would be really helpful, at least for me.” Physics culture usually rewards geniuses and creates intimidation for those who don’t identify as geniuses, so dismantling this notion such that students feel supported and safe enough to go to office hours and ask questions they do not know answers to would be very helpful.
Another aspect Ruby explains is important is for instructors to be approachable to students and clarify that students do not need to be experts to go to office hours: “Well, I think it’s so important for professors to explain that they are approachable and…also to get rid of this notion that students are supposed to come in already knowing everything that they want to ask, already knowing everything that they’re asking about. So I think that that’s on the professors to make abundantly clear that it’s okay, if you don’t know this, it’s okay, if you have a question, even if you think that this question might not be a good question there’s no such thing as a bad question. I think making that clear is so important…” In her opinion, instructors should have that role in making students feel safe and comfortable to attend office hours as well as be approachable enough.
@@
Encouraging peer collaboration: Another suggestion Ruby provides is in regard to collaboration with peers, she wishes that there were more opportunities to do so: “I wish there was more of a focus on group work, as well as individual work, because I think, I love to be able to talk with other students in a classroom as well as outside the classroom and learn from them and I don’t think there’s enough of that. And I think that that would also set us up a lot better for once we enter the real world and we do work with other people, so I think having an environment for that would be really, really great.” Not only would this be beneficial in introducing students to each other and forming strong connections that could support them throughout their program, but it would teach students how to collaborate with each other, as Ruby mentions.
@@
Restructuring in-class groups: Chloe, who is an introvert and has difficulty breaking barriers to collaborate with peers, suggests that instructors should change how groups are formed when it comes to in-class groups, such as: “assigning people to different groups and having rotations where people aren’t always stuck with the same group but also have enough time working with other people to break down these barriers, to force people to become friends…” Rather than keeping the same groups throughout the semester, rotating groups would help students meet each other. “[if] teachers have control over what the format of the classroom is and they can decide if it will just be lectures, or if they will work with groups inside the classrooms.” She justifies this by saying that instructors have control over how they structure their classes, so they can choose to incorporate group work in their classrooms and decide how groups are formed.
@@
Paulina also says that being able to communicate with her instructors about whether or not she feels comfortable with her partner would be important: “if we’d also have the choice to be like Hey, I don’t necessarily feel comfortable being partnered with this person, and then you wouldn’t have to. It would just be a conversation that you can have [with] the professor and then that kind of also alleviates the stress of engaging with someone who perhaps you haven’t had a positive experience with before.” It seems from Paulina negative experiences, a lot of the onus is on the student to bring this issue up. However, Paulina suggests that the instructor should make it clear that students can switch partners or provide feedback especially if they do not have good experiences with their partners. This might make it easier for students to switch partners.
@@
Instructors reinforcing disciplinary actions: Paulina emphasizes that instructors have responsibility in addressing inequities in their classroom. Specifically she refers to how students interact with each other. She also mentions that instructors have to be aware of any issues in their classrooms in order to properly intervene and take action: “It has to start off with the professor, I guess, because since the professor is the one that’s kind of in charge of engaging the students with one another, if they’re not aware of it, the problem where women in STEM don’t feel, women of color don’t feel as if they have an accurate voice in the conversations being had and there’s no way for them to personally intervene and frame their classroom to encourage the voices of women of color.” Again, we see a student emphasize that instructors have to be the ones who enact change. Paulina says that to do so, they have to be aware of what is going on in their classrooms, and with their students, otherwise how could they create change? She also says that they have to be aware of the issues in their classes in order to accurately discuss them.
@@
Check-in with students: Paulina shares that it would be beneficial for instructors to check-in with students: “I know it’s difficult for some people to have conversations in a large setting, so having one on one meetings, asking students how they feel in the space and [whether] they feel comfortable, they feel their questions are being answered, not just from the professor, but what’s their interactions with students and that’s also really helpful, because some students’ fear personally going up to professor and speaking up, just because you don’t necessarily know how it’s going to be met, and what kind of response they’re going to give you.” These types of conversations would help instructors understand how their students experience their classrooms and what their interactions with peers are like.
@@
Workshops to address biases against women and women of color: Following her suggestion for instructors to take action in their classrooms, Paulina also suggests that having workshops involving both instructors and students would be useful to address big issues: “I think, also just having interactive workshops between students and professors and maybe other faculty, where there is proper amount of attention given [to] this subject. Then, that’s when change can happen, I think it has to be something that is continuous and is made an apparently big deal because if it’s just kind of like something under the rug or it’s again only discussed once or twice, maybe in passing, then, most people would just not really take it to heart, it has to come to include the engagement of professors…and students, it can’t just be one or the other, because then they can’t really be resources for the students in the classroom.” Here she calls for workshops to be a space where issues in physics can be highlighted as current issues for the department to work on, rather than having them be mentioned once and then them being forgotten or not addressed.
@@
Frequent conversations with students about inequities: Paulina recalls some advice from someone she met about their instructor taking time out of class to have students engage with each other. That person conveyed to her that: “instead of focusing on solving like physics problems, we would engage with each other, and we would kind of provide a space where students can share their experiences if they wanted to and if no one felt really comfortable doing so, we were just kind of having conversations where I guess, students and professors would engage with the fact that, there is a lack of equity in the classroom for students of color and so having these, I guess, in-class workshops or in-class discussions, made it…[feel] as if it was important or situation that required a class time, for example. And so, people could engage with either other students of color or professors, or like dismantle any kind of stereotypes, or ideas that White students might have about the experiences of women or women of color in the classroom.” Paulina shares that having conversations frequently is useful because: “not only [do] the students begin to expect these kinds of conversations, but as time went on, you can see that there is more openness to ask more questions all around the classroom…” So the practice of normalizing important conversations might help other students feel comfortable asking for advice or insights. For example, Paulina says that attending a PWI, there are many White students who might not be aware that “ their thoughts or perceptions of students of colors are just biased or harmful to our learning experiences and so having a chance for us to have these continuous conversations allowed for them to come into the next workshop for example being Okay, well, I learned this from the last workshop, how can I expand on this, or how can I help other conversations and ask questions if I don’t really know how my actions or thoughts affect the students of color in the classroom.” She said that these workshops can provide an opportunity for people from dominant groups, such as White people, to be vulnerable about not understanding certain experiences and wanting feedback.
IV Comparison to Previous Studies

### v2

@@ IV.5.1 People providing academic support
In the previous themes, such as Family providing academic support, we discussed how some of the women already highlighted how their family provided support by helping them with their homework. For many of the women, their family members were also role models, especially if they had careers in science. For example, Cathy shares that her sister and her parents were some of her role models and provided foundations and inspiration for going into a science major. Many women noted that their professors were their role models. Some of these professors were also their academic or research advisors, thus serving multiple roles and having multiple contexts for providing support.
Another important source of support Chloe mentions is mentoring. The mentoring program she describes is part of a gender inclusive STEM club where upperclassmen partner with lowerclassmen. Chloe specifies that mentoring program is all girls and the upperclassmen help the lowerclassmen with their classes. She says that her mentor was: “really helpful and show[ed] me that I can apply for these research and these internships and to pursue these opportunities, yeah and just seeing people who are older and are currently doing what I am aspiring to do in the future is very inspiring.” This mentoring program connected her with more experienced students and particularly her mentor who guided her through opportunities such as research programs and internships.
IV.5.2 People providing emotional support
@@
Chloe shares that one of her professors is really inspiring because of a shared identity. Chloe considers herself to be an introvert and her professor is also quiet, so she feels inspired by her because: “she is kind of a quieter person and most of the professors that I’ve met, you ask them a question and then they talk very confidently for hours and she’s not like that…so meeting somebody who’s not like that but is still very successful in the field, yeah, that is very inspiring.” This type of role model is valuable as the stereotypical image of a physicist is usually either social and outgoing (like those portrayed in the media) or awkward and standoffish. Seeing someone who shares similar characteristics is helpful for someone like Chloe, who does not have the stereotypical personality of a physicist, to see herself as capable of being successful as one.
@@
IV.6 Suggestions to improve experiences
@@
In this section, we present the suggestions women provided to help improve their experiences. We separated these suggestions into two categories, the first is personal advice that has helped the women persist in physics and the second is suggestions for instructors they think would help improve their experiences.
IV.6.1 Personal Advice: The ”Band-Aid” Approach
@@
We note that although this personal advice may have helped a few of the women, it may not be useful for others. We also labeled this personal advice as The ”Band-Aid” Approach” because although this may be genuine advice from the women in physics, it only scratches the surface of the issues that traditionally marginalized students face within physics environments. In particular, this advice is more oriented towards personal action as opposed to more systemic change in the physics culture, which is discussed in a later sub-theme.
@@
Cathy has served on an undergraduate committee for physics majors in her program. She says that this opportunity has been beneficial for her because “…you can get involved in making whatever changes you feel would be helpful and having the department listen to the students…” Through her role on this committee, she was also able to host events for her peers, such as a physics GRE program and study program. She also emphasizes that: “I liked that once I was involved, I could then kind of make the events that I thought would be helpful so I’m glad I did that.” It appears that this may be personal advice based specifically upon the experience with undergraduate major committee that helped Cathy. We note that not every student may have this opportunity or may feel safe being in these types of roles. However, being involved in different types of departmental committees and activities based upon interest may be valuable for students.
@@
Ruby provides some advice that has helped her persist, which includes having a growth mindset. By adopting a growth mindset, she learned that the focus is more so on the process: “it’s about just practicing and trying and also just learning as much as you can. And I think that advice has really helped me in that feeling support, even when I’m not going to be even close to the smartest person in that room. [It] has really helped me.” Having growth mindset may in fact be useful for students to adopt. However, it is especially effective when instructors and the physics culture encourages and reflects a growth mindset as well so that the onus is not only on the student, like Ruby, to adopt a growth mindset in order to excel in physics. In particular, when instructors have growth mindset about their students’ potential and believe that all students can do well in physics, it can help students develop growth mindset and believe that they can become good at physics by working hard, working smart and taking advantage of resources.
@@
Francis shares that she thinks that believing in yourself would help future women in physics. She also encourages others to ignore people if they act like know-it-alls: “just because other people, especially men, are talking more and acting like they know more, it’s not necessarily true and you know, [you] are just as intelligent as them.” Thus, Francis’ advice based upon her personal experiences is to “believe in yourself” and although this is a good thing to do, it may be difficult for women in physics if they do not have the support and it does not appear to them that other people believe in them (as in the case of Paulina). This advice may be viewed by some as putting onus on the student to change to fit in to survive and make it through the physics culture. Having instructors believe in their students is necessary for traditionally marginalized students to truly believe in themselves with regard to excelling in physics. These positive changes can be heralded with changing the actual culture of physics that can be usually intimidating for many students.
IV.6.2 Advice for instructors
@@
Avoid forming opinions about students based on background: Karina shares some insight on how instructors should not make assumptions about students based on their background. This can be problematic because “…especially [at] a small school like [my college], where you might have professors for multiple years, you might come back to that professor that you’ve had your first semester and I think that’s important, that’s something that’s important for professors to know.” Besides potentially having the same instructor for multiple courses, these instructors also serve as research advisors and mentors, so it is especially important to have good perceptions of (and relationships with) students. In addition, having preconceived notions about students may unintentionally lead to instructors stereotyping students or making generalizations, particularly about their ability.
@@
Address and dismantle biases against women: Ruby shares that it is important for instructors to explicitly communicate in their actions and words that women are just as capable as men in terms of their physics ability. She says that instructors should “mak[e] it a very, very clear fact that…men are no better at physics than women. Just because there are more men in the history of physics does not mean that men in your classroom are going to be better at it than you are because of your gender, because that has no influence on how capable you are [of] learning the subject, even if people tell you otherwise.” Changing the physics culture to reflect beliefs about all students as equally capable of excelling in physics is key. Having instructors emphasize this and continue to do so not only through verbal affirmations but via their actions as well would be necessary to reflect how deeply they support this tenet.
@@
In one of her reflections regarding her gender identity, she provides a suggestion for instructors to address biases against women. For context, she emphasizes that she has to work hard to prove that she is capable: “I feel like as a woman, I have to prove myself for people to look at me and go, Oh, you’re smart and dedicated and you like science, I have to prove that first. It’s not implicitly assumed and nobody assumes that when they first meet me… for many men in my classes it’s the opposite, where it’s like you initially assume that this person is really smart and gets really good grades just because they’re male in science.” She believes increasing the number of women in physics programs could help reduce these biases: “if many other girls are also doing really well, I think it’s a self perpetuating cycle that as more women get involved in science, the more everybody… realizes that we’re on the same playing field.” So having instructors dismantle these stereotypes and biases about who belongs in physics and can excel in it through their actions can be very impactful.
@@
Reassure about lack of representation of women in physics: Ruby shares that she is aware of the historical lack of representation of women in physics but she would appreciate her department to acknowledge that while uplifting women: “I haven’t really ever been told …just because there aren’t a lot of women that did it before you, doesn’t mean you can’t do it, but it’d be so reassuring and helpful to hear…I think that that would be beneficial to sort of creating that comfortable environment and creating that confidence.” Thus, Ruby emphasizes that acknowledging this part of physics history would not only educate departmental members but also students would be aware of this going forward in their major. Following this acknowledgement, reassurance would be helpful so that women are not worried about majoring in physics and the department feels supportive to them.
@@
Create a better representation of women in physics: Esme shares that having more representation of women in physics would be beneficial: “seeing more women be the forefront, be the face of physics, because usually when we think of physics, we think of the men… but you know, seeing more women just in their discoveries is something that probably can be helpful.” Based upon Esme’s suggestion, highlighting more women physicists such as in lectures, seminars, newsletters, etc. can provide support for women. This can show them the achievements of women physicists who have made major contributions to physics historically and in the present, which can potentially improve their sense of belonging and persistence in physics.
@@
Provide reassurance during struggle: Ruby mentions how reassurance would be helpful especially when she does not feel confident in herself. For example, not doing well on an exam would be an opportunity to reassure students because: “Sometimes it’s scary and sometimes it makes you feel like am I in this and not able to do this?, and a lot of the times, my confidence is definitely shaken up by a test grade or something along those lines…” She also says that instructors should create a safe environment to get problems wrong because: “it’s okay to not always get the right answer…and that doesn’t mean you’re stupid and that doesn’t mean that you shouldn’t do what you’re passionate about…” Thus reframing mindsets about wrong answers as an opportunity to learn would be useful in addition to instructors reassuring students that they do belong in physics and that it is okay to not do well on an exam.
@@
Improve office hours: Ruby also provides several suggestions for instructors to improve their classrooms, one of them is to improve the perception and purpose of office hours. She says, “making a more accepting environment and comfortable environment to make mistakes and to not always know the answer… so that people can feel confident, even when they don’t know everything.” She also calls for instructors to dismantle the idea that students should be knowledgeable about the questions they ask. She says, “that’s on the professors to make abundantly clear that it’s okay, if you don’t know this, it’s okay, if you have a question, even if you think that this question might not be a good question, there’s no such thing as a bad question.” Thus, the responsibility is on the instructor to create this environment that feels safe to ask questions. According to Ruby, this is a problem with the physics culture, that “ people are terrified to even try to answer a question in the fear of getting it wrong…” Physics culture usually rewards geniuses and creates intimidation for those who don’t identify as geniuses, so dismantling this notion such that students feel supported and safe enough to go to office hours and ask questions they do not know answers to would be very helpful. Ruby also adds that instructors should be approachable to students so that students are not as afraid to ask them questions.
@@
Encourage peer collaboration: Another suggestion Ruby provides is in regard to collaboration with peers. She wishes that there were more opportunities to work with peers because: “that would also set us up a lot better for once we enter the real world and we do work with other people, so I think having an environment for that would be really, really great.” Not only would this be beneficial in introducing students to each other and forming strong connections that could support them throughout their program, but it would teach students how to collaborate with each other, as Ruby mentions.
@@
Restructure in-class groups: Chloe, who is an introvert and has difficulty breaking barriers to collaborate with peers, suggests that instructors should change how groups are formed when it comes to in-class groups, such as: “assigning people to different groups and having rotations where people aren’t always stuck with the same group but also have enough time working with other people to break down these barriers, to force people to become friends…” Rather than keeping the same groups throughout the semester, rotating groups would help students meet each other. “[if] teachers have control over what the format of the classroom is and they can decide if it will just be lectures, or if they will work with groups inside the classrooms.” She justifies this by saying that instructors have control over how they structure their classes, so they can choose to incorporate group work in their classrooms and decide how groups are formed.
@@
Paulina also says that being able to communicate with her instructors about whether or not she feels comfortable with her partner would be important. She explains, “It would just be a conversation that you can have [with] the professor and then that kind of also alleviates the stress of engaging with someone who perhaps you haven’t had a positive experience with before.” It seems from Paulina negative experiences, a lot of the onus is on the student to bring this issue up. However, Paulina suggests that the instructor should make it clear that students can switch partners or provide feedback especially if they do not have good experiences with their partners. This might make it easier for students to switch partners.
@@
Reinforce disciplinary actions: Paulina emphasizes that instructors have responsibility in addressing inequities in their classroom. Specifically she refers to how students interact with each other. She also mentions that instructors have to be aware of any issues in their classrooms in order to properly intervene and take action: “It has to start off with the professor, I guess, because since the professor is the one that’s kind of in charge of engaging the students with one another, if they’re not aware of it, the problem where women in STEM don’t feel, women of color don’t feel as if they have an accurate voice in the conversations being had and there’s no way for them to personally intervene and frame their classroom to encourage the voices of women of color.” Again, we see a student emphasize that instructors have to be the ones who enact change. Paulina says that to do so, they have to be aware of what is going on in their classrooms, and with their students, otherwise how could they create change? She also says that they have to be aware of the issues in their classes in order to accurately discuss them.
@@
Check-in with students: Paulina shares that it would be beneficial for instructors to check-in with students especially if some students struggle with large group settings. She describes what these check-ins would look like: “having one on one meetings, asking students how they feel in the space and [whether] they feel comfortable, they feel their questions are being answered, not just from the professor, but what’s their interactions with students?” Paulina also adds that this can create a better relationships between the instructor and students. These types of conversations would help instructors understand how their students experience their classrooms and what their interactions with peers are like.
@@
Engage in conversations with students about inequities: Following her suggestion for instructors to take action in their classrooms, Paulina also suggests that having workshops involving both instructors and students would be useful to address big issues. According to her, this would bring a “proper amount of attention” to these issues in physics. She highlights that “that’s when change can happen, I think it has to be something that is continuous and is made an apparently big deal because if it’s just kind of like something under the rug or it’s again only discussed once or twice, maybe in passing, then, most people would just not really take it to heart…” Here she calls for workshops to be a space where issues in physics can be highlighted as current issues for the department to work on, rather than having them be mentioned once and then them being forgotten or not addressed.
@@
Paulina also recalls some advice from someone she met about their instructor taking time out of class to have students engage with each other. That person conveyed to her that rather than focus on solving physics problems, students “engage with each other, and we would kind of provide a space where students can share their experiences if they wanted to… students and professors would engage with the fact that there is a lack of equity in the classroom for students of color and so having these, I guess, in-class workshops or in-class discussions, made it…[feel] as if it was important or situation that required a class time…” Paulina shares that having conversations frequently is useful because: “not only [do] the students begin to expect these kinds of conversations, but as time went on, you can see that there is more openness to ask more questions all around the classroom…” So the practice of normalizing important conversations might help other students feel comfortable asking for advice or insights. For example, Paulina says that attending a PWI, there are many White students who might not be aware that they have biases or harmful perceptions of students of color that is “harmful to our learning experiences and so having a chance for us to have these continuous conversations allowed for them to come into the next workshop for example being Okay, well, I learned this from the last workshop, how can I expand on this, or how can I help other conversations and ask questions if I don’t really know how my actions or thoughts affect the students of color in the classroom.” She said that these workshops can provide an opportunity for people from dominant groups, such as White people, to be vulnerable about not understanding certain experiences and wanting feedback.
V Broader Discussion and Brief Comparison to Previous Studies
We now situate this work by summarizing how it compares to previous qualitative studies related to experiences of undergraduate women majoring in physics. We start with brief summaries of the two studies conducted by Johnson (2020) and Santana and Singh (2023).
V.1 Johnson’s 2020 study
In 2020, Johnson conducted open-ended interviews in a small physics department at a small liberal arts college in the US with 6 undergraduate women majoring in physics, and a focus group consisting of third year physics students. She asked undergraduate women, during the individual interviews, questions like “Tell me your life story in physics.” Four physics faculty in the department were also interviewed but we will not focus on that here. Out of the individual interviews with undergraduate women, three were women of color and three were White women. At this institution, 25% of physics majors are women. For focus groups, she asked students about their trajectory in physics, what students like about their physics classes, what could be improved, and what ideas they might have for attracting more women to physics.
We chose to compare the study discussed here in a medium-size physics department at a small liberal arts college with this study because of the overwhelmingly positive physics culture that both students and faculty talked about [70]. We believe this physics department can serve as a good model for an equitable and inclusive physics environment where all students thrive and feel safe [70].
V.2 Santana and Singh’s 2023 study
Santana and Singh conducted semi-structured, empathetic interviews with 16 undergraduate women physics and astronomy majors at a large research university in the US. At the time when the interviews were conducted, this number of women interviewed represents approximately 60% of the women students in the physics and astronomy department (23% women). Six out of 16 interviews were analyzed which included a wide range of experiences. The goals and the protocols were similar as this current study.
Findings from this study revealed a masculine physics culture that did not support undergraduate women [23]. In particular, this study revealed that women had many negative interactions with their male peers and male instructors. the findings point to an inequitable and non-inclusive physics learning environment unlike the one described in Johnson’s study. One reason we chose this study for this comparative analysis is that it highlights undergraduate women’s experiences in a physics department at a large research university in the US that may represent a more prototypical physics department.
V.3 Comparing the three studies